# Masked Autoencoders are PDE Learners

### v2

@@ 4.1 Equations Considered
The KdV-Burgers equation (KdV-B) contains the Heat, Burgers, KdV equations as corner cases modulated by coefficients ($\alpha,\beta,\gamma$) [8; 58]. Furthermore, periodic boundary conditions are used with a forcing function defined by $\delta(x,t)$. For 1D multi-resolution experiments, PDE data from the KdV-Burgers equation is downsampled to variable spatial resolutions.
@@
Although similar to the 1D KdV-Burgers equation, the 1D Heat and inviscid Burgers [8] equation are not explicitly seen during pretraining and are used to evaluate generalization to PDEs resembling pretraining samples. Furthermore, to evaluate extrapolation to unseen boundary conditions (BCs), samples of the Heat equation are also generated with Dirichlet and Neumann BCs in addition to periodic BCs.
@@
The Advection (Adv), Wave, and parameter-dependent Kuramoto-Sivashinsky (KS) [59] equations are considered to evaluate downstream performance to completely new equations; the equations contain PDE terms that are unseen during pretraining. Additionally, the Wave equation is generated with Dirichlet and Neumann BCs [8] to evaluate unseen BCs on novel PDE dynamics.
@@
The 2D Heat, Advection (Adv), and scalar Burgers [60] equations are considered for both pretraining and downstream evaluation. For 2D multi-resolution experiments, data from these equations are downsampled to variable spatial resolutions.
@@
Following the setup from Li et al. (2020), we consider the incompressible Navier-Stokes (NS) equations in vorticity form, but randomly sample the viscosity $\nu$ and forcing function $f(x)$ amplitude. To ensure consistency with the pretraining dataset, our experiments model NS dynamics as a scalar vorticity field; from this the velocity field can be derived.
@@
Figure 2: Example results after training on the 1D KdV-Burgers equation only with a masking ratio of 75%. For each triplet, we show the masked PDE (left), the MAE reconstruction (middle), and the ground-truth (right), and plot space and time on the $x$ and $y$ axes respectively. The MAE is able to reconstruct multiple resolutions of KdV-Burgers data, as well as extrapolate to the 1D Heat and inviscid Burgers equations, which are unseen during training. For the 1D Advection and KS equations, which contain novel PDE terms ($u_{x},u_{xxxx}$), the extrapolation performance is limited.
@@ 4.2 MAE Pretraining
MAE models are trained on 10000 samples of 1D KdV-Burgers PDE data in 1D and 12288 samples of 2D Heat, Advection, and Burgers PDE data in 2D. We display example results from masked pretraining in Figures 3 and 3. A notable difference from vision and video domains is that physics does not follow human-recognizable structure or descriptions (e.g. backgrounds, actions, faces, shapes, etc.); furthermore, in addition to the overall meaning, the numerical accuracy of the reconstruction is important. Despite this, MAE models are able to capture underlying physics and reconstruct PDEs within the pretraining set well—both in 1D and 2D, and at high masking ratios (75% and 90%). In general, for PDEs that are similar to those seen during pretraining, such as the 1D Heat or inviscid Burgers equation, MAE models tend to extrapolate well. Furthermore, given information about the spatial discretization, MAE models can adapt to different PDE resolutions when trained to reconstruct multi-resolution inputs. This is true in 2D as well, with example results shown in Appendix C.2.
For PDEs that contain novel equation terms or boundary conditions (BCs), the MAE extrapolation performance is limited. Zero-shot reconstruction of the 1D Advection and KS equations shows mild trends, while reconstruction of the 2D Navier-Stokes equations is ineffective. To address this gap and investigate whether MAE models can perform on complex high-resolution physics with multiple variables, we train an MAE model to reconstruct pressure and velocity on 2D smoke buoyancy data with varying buoyancy factors [18] and qualitatively show its reconstruction in Appendix C.3. In general, MAE models can adapt to complex scenarios and multiple physical variables; however, many of the fine details (e.g. eddies, shedding) become lost at high masking ratios.
Since the 1D KdV-Burgers equation is solved with periodic BCs, we evaluate MAE extrapolation to Dirichlet and Neumann BCs when reconstructing the 1D Heat and Wave equations, shown in Appendix C.1. Overall trends remain consistent; the Heat equation, being more similar to the KdV-Burgers equation, shows limited reconstruction performance when extrapolating to novel BCs, while the Wave equation introduces a novel PDE term ($u_{tt}$) and initial condition (Gaussian pulse), and as a result the zero-shot reconstruction is poor.
Figure 4: MAE embeddings of various PDEs after masked pretraining; the MAE latent space shows structure despite not seeing labels. A: 1D KdV-Burgers equation, colored by $\alpha$. B: 1D Advection equation, colored by $c$. C: 1D Heat, Burgers, Advection, and KS equations, colored by PDE. D: 1D KdV-Burgers equation, colored by resolution. E: 1D Heat equation, colored by boundary condition. F: 2D Heat, Advection and Burgers equations, colored by $\nu$ and $c$.
@@ 4.3 Latent Space Evaluation
Latent Embeddings
To better understand the latent representation learned by masked pretraining, we use the MAE encoder to embed various PDE validation samples and visualize embeddings with t-SNE [61] in Figure 4. Through self-supervised learning, MAE models can learn trends in PDEs without labeled data and in equations outside the training set. For example, through masked reconstruction of the 1D KdV-Burgers equation, MAE models can learn coefficient-dependent trends (Figure 4A). This can be applied to PDEs not seen during training, as the same MAE model can distinguish samples from the Advection equation based on wave speed $c$ (Figure 4B). This extrapolation is also observed when embedding samples from different PDEs; representations learned from pretraining on the KdV-Burgers equations allow models to cluster PDE samples from the Heat, Burgers, Advection, and KS equations (Figure 4C). Lastly, MAE models are able to distinguish varying resolutions of PDE data after multi-resolution training, suggesting high model capacity across both diverse PDEs and discretizations (Figure 4D).
@@
Latent PDE Arithmetic
MAE encoders have shown strong capabilities in extracting information from self-supervised PDE learning, creating opportunities to operate in this latent space. After pretraining on the 1D KdV-Burgers equation, we consider an arithmetic task where samples of the Heat and Burgers equation are embedded and added in the latent space before being decoded to a novel solution (Figure 5). Concretely, we generate 1D Heat data from the PDE: $\partial_{t}u-\nu\partial_{xx}u=0$,
Figure 6: Latent PDE interpolation
1D inviscid Burgers data from the PDE: $\partial_{t}u+u\partial_{x}u=0$, and 1D Heat-Burgers data from adding the two PDEs: $\partial_{t}u-\nu\partial_{xx}u+u\partial_{x}u=0$. When given identical initial conditions and coefficients, PDE reconstructions obtained from summing latent Heat and Burgers embeddings qualitatively resemble solutions of the combined Heat-Burgers PDE. Notably, the MAE encoder/decoder has never seen the 1D Heat or Burgers equation on their own.
In addition, if different latent embeddings can be added, the weighting of each embedding can be varied to control the resemblance of the reconstruction. We empirically show this in Figure 6 for the Heat-Burgers case presented; however, we anticipate that this work can be expanded and improved on. In particular, the decoder has only seen masked data during training, and as a result latent reconstructions must be done with partially encoded information.
@@
4.4 PDE Feature Prediction
To evaluate the latent representations learned from masked pretraining, we regress PDE coefficients and classify PDE boundary conditions, equation families, and spatial resolutions from an initial time window. Regression tasks are separated by PDE (KdV-B, KS, Heat, Adv, Burgers, NS). Further, classification tasks are formulated as: (HeatBC): predicting Periodic, Dirichlet, or Neumann BCs from the Heat equation, (WaveBC): predicting Dirichlet or Neumann BCs from the Wave equation, (PDEs): predicting unseen PDEs from Heat, Adv, Burgers, and KS equation samples, and (Res.): predicting resolutions from $n_{x}=\{50,60,70,80,90,100\}$ in 1D or $(n_{x},n_{y})=\{(48,48),(52,52),(56,56),(60,60),(64,64)\}$ in 2D. Lastly, several model variants are evaluated: a randomly initialized ViT baseline (Base), a pretrained, frozen MAE encoder with a linear head (MAEf), and a pretrained, fine-tuned MAE encoder (MAE). Results are shown in Tables 2 & 2, with full error bars in Appendix E.1.
In 1D, the frozen MAEf encoder is able to outperform a supervised baseline on the KdV-B, Heat, PDEs, and Res. tasks despite never seeing labels throughout training. Further performance gains can be realized by allowing the MAE encoder to fine-tune on labeled data, and can outperform random initialization on unseen equations and BCs. An exception to this is the Advection and Wave equations. We hypothesize that these PDEs are heavily governed by the wave speed $c$ and boundary conditions (bouncing vs. reflecting waves) and are simple trends that supervised models can learn quickly.
Table 3: Conditional 1D PDE time-stepping and super-resolution. Models are trained on 2000 held-out fine-tuning samples to predict or upsample physics across several settings. Validation errors are reported as normalized L2 loss (time-stepping) or RMSE×$\times 10^{-1}$ (SR) summed over all timesteps and averaged across five seeds.

### v3

@@ 4.1 PDEs and Datasets
The KdV-Burgers equation (KdV-B) contains the Heat, Burgers, KdV equations as corner cases modulated by coefficients ($\alpha,\beta,\gamma$) (Brandstetter et al., 2022a; Jeffrey and Mohamad, 1991):
@@
Both initial conditions and forcing terms are generated from the periodic function $\delta(t,x)$, where we uniformly sample $A_{j}\in[-0.5,0.5],\omega_{j}\in[-0.4,0.4],l_{j}\in\{1,2,3\},\phi_{j}\in[0,2\pi)$ while fixing $J=5,L=16$.
To generate diverse PDE samples, the coefficients are sampled uniformly from $\alpha\in[0,3],\beta\in[0,0.4],\gamma\in[0,1]$. Furthermore, samples are generated with a discretization $(n_{t},n_{x})=(250,100)$ on an interval $x=[0,16]$ from $t=0$ to $t=2$. For 1D multi-resolution experiments, PDE data from the KdV-Burgers equation is downsampled to variable spatial resolutions.
@@
The 1D Heat and inviscid Burgers (Brandstetter et al., 2022a) equations are used to evaluate performance on PDEs that are a subset of pretraining samples. Furthermore, to evaluate extrapolation to unseen boundary conditions (BCs), samples of the Heat equation are also generated with Dirichlet and Neumann BCs in addition to periodic BCs.
@@
We solve the equations with the same periodic BCs, initial conditions, and forcing function setup and as the 1D KdV-Burgers equation, but by setting the appropriate coefficient values. Specifically, we uniformly sample $\nu=\beta\in[0.1,0.8]$ for the Heat equation and fix $\alpha=0.5,\beta=0$ to model the inviscid Burgers equation. These equations are also solved with a discretization $(n_{t},n_{x})=(250,100)$ on an interval $x=[0,16]$ from $t=0$ to $t=2$. To generate data for the Heat equation that enforces Dirichlet or Neumann boundary conditions, we write the Heat equation in its variational form after an implicit Euler discretization:
[…]
This formulation can be solved using FEniCS (Alnaes et al., 2015; Logg et al., 2012). To simplify the boundary value problem, we set the forcing term $f=0$. Furthermore, we set [$u(x=0)=u(x=L)=0$] to enforce Dirichlet BCs, and [$\partial_{x}u(x=0)=\partial_{x}u(x=L)=0$] to enforce Neumann BCs; however, in both of these cases, the initial conditions in Equation 1 need to be modified to respect the new BCs.
@@
The Advection (Adv), Wave, and parameter-dependent Kuramoto-Sivashinsky (KS) (Lippe et al., 2023) equations are considered to evaluate downstream performance to new equations; the equations contain PDE terms that are unseen during pretraining. Additionally, the Wave equation is generated with Dirichlet and Neumann BCs (Brandstetter et al., 2022a) to evaluate unseen BCs on novel PDE dynamics.
@@
For the 1D Advection equation, initial conditions are generated according to Equation 1, the wave speed is uniformly sampled from $c\in[0.1,5]$, and periodic BCs are used. The solution domain and discretization are the same as previous cases, with $(n_{t},n_{x})=(250,100)$, $x=[0,16]$, and time ranging from $t=0$ to $t=2$.
For the 1D Wave equation, we solve with Dirichlet ($u(x=0)=u(x=L)=0$) and Neumann ($\partial_{x}u(x=0)=\partial_{x}u(x=L)=0$) BCs, resulting in waves that either bounce or reflect off boundaries. The wave speed is fixed at $c=2$, and the initial condition is a Gaussian pulse with unit amplitude and with its peak randomly sampled on the spatial domain. Lastly, the equation is solved from $t=0$ to $t=100$ on the interval $x=[-8,8]$ with a discretization $(n_{t},n_{x})=(250,100)$.
For the 1D KS equation, we use periodic BCs with initial conditions from Equation 1. Following the data setup proposed by Lippe et al. (2023), we additionally uniformly sample $\nu\in[0.75,1.25]$ to vary the second-order term in the KS equation. Furthermore, due to the unique dynamics of the KS equation, we solve the PDE from $t=0$ to $t=100$ on the interval $x=[0,64]$ with a discretization of $(n_{t},n_{x})=(100,100)$.
@@
The 2D Heat, Advection (Adv), and scalar Burgers (Rosofsky et al., 2023) equations are considered for both pretraining and downstream evaluation. For 2D multi-resolution experiments, data from these equations are downsampled to variable spatial resolutions.
@@
Following the setup from Li et al. (2020), we consider the incompressible Navier-Stokes (NS) equations in vorticity form, but randomly sample the viscosity $\nu$ and forcing function $f(x)$ amplitude. To ensure consistency with the pretraining dataset, our experiments model NS dynamics as a scalar vorticity field; from this the velocity field can be derived from the Biot-Savart Law.
@@
The solution is solved on a grid $(n_{t},n_{x},n_{y})=(100,64,64)$ on a solution domain $(x,y)=[0,1]^{2}$ from $t=0$ to $t=25$. PDE parameters are uniformly sampled from $\nu\in\{\{1,2,3,4,5,6,7,8,9\}\times 10^{-\{6,7,8,9\}}\}$ and $A\in\{1,2,3,4,5,6,7,8,9,10\}\times 10^{-3}$. Lastly, initial conditions $\omega_{0}$ are sampled according to Li et al. (2020) from a Gaussian random field.
4.2 Data Augmentations
We implement Lie Point Symmetry Data Augmentations according to Brandstetter et al. (2022b), including shifting and resampling PDE solutions with the Fourier shift theorem. Since we only augment PDE samples during pretraining, we consider symmetry groups for the 1D KdV-Burgers equation, as well as the 2D Heat, Advection, and Burgers equations. The 1D KdV-Burgers equation has the following Lie subalgebras (Ibragimov, 1993):
Taking the exponential map results in the following Lie groups (Ibragimov, 1993):
[…]
For the 2D Heat, Advection, and Burgers equations, there are many possible Lie subalgebras (Ibragimov, 1993). For simplicity, we only consider a basic subset of these that apply to all three equations, however, there is ample room to implement more symmetries:
These result in the following Lie groups in 2D:
[…]
5 Results
In this section, we provide the results of different experiments designed to understand the capabilities of masked autoencoders. Firstly, we would like to understand the reconstruction capabilities of the masked autoencoder, or if the pretraining goal given by the masked objective is being met. Once MAE performance is validated on its pretraining objective, we seek to understand and visualize the representations learned by the MAE during pretraining. This can be done by projecting the latent embeddings to a lower dimension to visualize qualitative trends. Although insightful, another more rigorous evaluation is the regression/classification of physical variables. These are easy quantities to derive and are usually trivially known a priori, however, they can serve as a probe to quantitatively gauge model knowledge rather than rely on qualitative latent trends. This is analogous to a linear probe used to predict image rotations or grayscale vs. color in self-supervised learning for computer vision (Chen et al., 2020); indeed, coefficient regression has been used in prior PDE literature to gauge model performance after self-supervised pretraining (Mialon et al., 2023).
While these three tasks (masked reconstruction, latent visualization, variable regression/classification) can provide knowledge, they are generally not useful on their own. To extend masked pretraining to practical tasks, we consider using a pretrained encoder to improve PDE time-stepping or super-resolution. In particular, we are interested in whether the representations learned during masked pretraining can help in diverse physics scenarios by providing additional context to neural PDE solvers during time-stepping or super-resolution.
Figure 2: Example results after training on the 1D KdV-Burgers equation with a masking ratio of 75%. For each triplet, we show the masked PDE (left), the MAE reconstruction (middle), and the ground-truth (right), and plot space and time on the $x$ and $y$ axes respectively. The MAE can reconstruct multiple resolutions of KdV-Burgers data and interpolate to the 1D Heat and inviscid Burgers equations. For the 1D Advection and KS equations, which contain novel PDE terms ($u_{x},u_{xxxx}$), the extrapolation performance is limited.
@@ 5.1 MAE Pretraining
MAE models are trained on 10000 samples of 1D KdV-Burgers PDE data in 1D and 12288 samples of 2D Heat, Advection, and Burgers PDE data in 2D. We display example results from masked pretraining in Figures 3 and 3. A notable difference from vision and video domains is that physics does not follow human-recognizable structure or descriptions (e.g. backgrounds, actions, faces, shapes, etc.); furthermore, in addition to the overall meaning, the numerical accuracy of the reconstruction is important. Despite this, MAE models are able to capture underlying physics and reconstruct PDEs within the pretraining set well—both in 1D and 2D, and at high masking ratios (75% and 90%). In general, for PDEs that are similar to those seen during pretraining, such as the 1D Heat or inviscid Burgers equation, MAE models tend to interpolate well. Furthermore, given information about the spatial discretization, MAE models can adapt to different PDE resolutions when trained to reconstruct multi-resolution inputs. This is true in 2D as well, with example results shown in Appendix B.2.
For PDEs that contain novel equation terms or boundary conditions (BCs), the MAE extrapolation performance is limited. Zero-shot reconstruction of the 1D Advection and KS equations shows mild trends, while reconstruction of the 2D Navier-Stokes equations is ineffective. To address this gap and investigate whether MAE models can perform on complex high-resolution physics with multiple variables, we train an MAE model to reconstruct pressure and velocity on 2D smoke buoyancy data with varying buoyancy factors (Gupta and Brandstetter, 2022) and qualitatively show its reconstruction in Appendix B.3. In general, MAE models can adapt to complex scenarios and multiple physical variables; however, many of the fine details (e.g. eddies, shedding) become lost at high masking ratios.
Since the 1D KdV-Burgers equation is solved with periodic BCs, we evaluate MAE extrapolation to Dirichlet and Neumann BCs when reconstructing the 1D Heat and Wave equations, shown in Appendix B.1. Overall trends remain consistent; the Heat equation, being more similar to the KdV-Burgers equation, shows limited reconstruction performance when extrapolating to novel BCs, while the Wave equation introduces a novel PDE term ($u_{tt}$) and initial condition (Gaussian pulse), and as a result the zero-shot reconstruction is poor.
@@ 5.2 Latent Space Evaluation
To better understand the latent representation learned by masked pretraining, we use the MAE encoder to embed various PDE validation samples and visualize embeddings with t-SNE (van der Maaten and Hinton, 2008) in Figure 4. Through self-supervised learning, MAE models can learn trends in PDEs without labeled data and with limited extrapolation abilities. For example, through masked reconstruction of the 1D KdV-Burgers equation, MAE models can learn coefficient-dependent trends (Figure 4A). This can be applied to PDEs not seen during training, as the same MAE model can distinguish samples from the Advection equation based on wave speed $c$ (Figure 4B). This extrapolation is also observed when embedding samples from different PDEs; representations learned from pretraining on the KdV-Burgers equations allow models to cluster PDE samples from the Heat, Burgers, Advection, and KS equations (Figure 4C). Lastly, MAE models are able to distinguish varying resolutions of PDE data after multi-resolution training, suggesting high model capacity across both diverse PDEs and discretizations (Figure 4D).
Figure 4: t-SNE embeddings of various PDEs. Plots show embeddings before and after using the MAE to encode samples, shown on the top and bottom. The MAE latent space shows structure despite not seeing labels of coefficients, PDEs, or BCs. A: 1D KdV-Burgers equation, colored by $\alpha$. B: 1D Advection equation, colored by $c$. C: 1D Heat, Burgers, Advection, and KS equations, colored by PDE. D: 1D KdV-Burgers equation, colored by resolution. E: 1D Heat equation, colored by boundary condition. F: 2D Heat, Advection and Burgers equations, colored by $\nu$ and $c$.
@@
5.3 PDE Feature Prediction
To evaluate the latent representations learned from masked pretraining, we regress PDE coefficients and classify PDE boundary conditions, equation families, and spatial resolutions from an initial time window. Regression tasks are separated by PDE (KdV-B, KS, Heat, Adv, Burgers, NS). Further, classification tasks are formulated as: (HeatBC): predicting Periodic, Dirichlet, or Neumann BCs from the Heat equation, (WaveBC): predicting Dirichlet or Neumann BCs from the Wave equation, (PDEs): predicting unseen PDEs from Heat, Adv, Burgers, and KS equation samples, and (Res.): predicting resolutions from $n_{x}=\{50,60,70,80,90,100\}$ in 1D or $(n_{x},n_{y})=\{(48,48),(52,52),(56,56),(60,60),(64,64)\}$ in 2D.
Several model variants are evaluated: a randomly initialized ViT baseline (MAEb), a pretrained, frozen MAE encoder with a linear head (MAEf), and a pretrained, fine-tuned MAE encoder (MAE). For these MAE models, regression and classification are performed using a CLS token and projecting the CLS embedding to the number of prediction features through a simple MLP. The baseline encoder is a ViT with the same model size and architecture as the pretrained MAE encoder, and when the MAE encoder is frozen, only the MLP head receives gradient updates. Additionally, we train MLP and CNN models to regress coefficients or classify PDE features in order to benchmark the difficulty of these tasks. Since these models cannot process variable sized inputs, the Res. experiment is not performed for these simpler baselines. Results are shown in Tables 4 & 4, with full error bars in Appendix E.1.
In 1D, the frozen MAEf encoder is able to outperform a supervised baseline on the KdV-B, Heat, PDEs, and Res. tasks despite never seeing labels throughout training. Further performance gains can be realized by allowing the MAE encoder to fine-tune on labeled data, and can outperform random initialization on unseen equations and BCs. An exception to this is the Advection and Wave equations. We hypothesize that these PDEs are heavily governed by the wave speed $c$ and boundary conditions (bouncing vs. reflecting waves) and are simple trends that supervised models can learn quickly. Indeed, the simple MLP and CNN baselines are able to achieve perfect classification accuracy on the Wave equation, likely since the wave either reflects or bounces off the boundaries according to the BC.
@@ 5.3 PDE Feature Prediction
In 2D, the effects of fine-tuning are more pronounced. Within the pretraining set, only in the 2D Burgers task does freezing the MAE encoder outperform a supervised baseline; nevertheless, masked pretraining serves as a good initialization for supervised fine-tuning. In addition, despite differing physics, prior knowledge from simpler 2D PDEs seems to benefit regression on the Navier-Stokes equations. When classifying 2D Heat, Advection, and Burgers data based on their discretization, MAE models greatly benefit from pretraining on multi-resolution data. We hypothesize that in 2D PDEs, variable spatial resolutions can be challenging to distinguish due to flattening the spatial dimension when patchifying inputs, whereas in 1D PDEs the data is already flattened.
5.4 Conditional Time-stepping and Super-resolution
1D Experiments
To evaluate the use of the MAE encoder for practical tasks, we train a neural solver on various 1D PDEs to predict or upsample physics. For prediction, or time-stepping, models are given solutions are time $t$ and queried to predict solutions at at time $t+1$. For upsampling, or super-resolution, models are given a low-resolution solution at time $t$ and queried to predict a high-fidelity solution at time $t$. For our experiments in 1D, we consider five PDEs (KdV-B, Heat, Burgers, Adv, KS) as well as two PDEs under varying boundary conditions (HeatBC, WaveBC) and predicting physics on various resolutions of the KdV-Burgers equation (Res.). Time-stepping is performed autoregressively by predicting multiple timesteps simultaneously to reduce error accumulation (Brandstetter et al., 2022a). Furthermore, the pushforward trick (Brandstetter et al., 2022a) is implemented. This adds model noise to inputs during training by making a prediction of a future timestep and using that prediction as a noised input the model; importantly gradients are not calculated for the initial pass. Lastly, we test on FNO (Li et al., 2020) and Unet architectures (Gupta and Brandstetter, 2022), (Ronneberger et al., 2015), and add conditioning information to hidden states after convolutions (Ho et al., 2020; Nichol and Dhariwal, 2021).
Table 5: Conditional 1D PDE time-stepping and super-resolution. Models are trained on 2000 held-out fine-tuning samples to predict or upsample physics across several settings. Validation errors are reported as normalized L2 loss (time-stepping) or RMSE×$\times 10^{-1}$ (SR) summed over timesteps and averaged across five seeds.